OpenStudy (anonymous):

Simplify: -7(3x + 2)

OpenStudy (anonymous):

You can get this one... distribute (multiply) the -7 across the parenthesis.

OpenStudy (anonymous):

-7X3x

OpenStudy (anonymous):

-7X2

OpenStudy (anonymous):

right.... -7(3x+2) = -7(3x) + (-7)(2) = -21x -14

OpenStudy (anonymous):

-35x?

OpenStudy (anonymous):

no, the "-14" isn't an x term... you can't add it to the -21x term

OpenStudy (anonymous):

oh.

OpenStudy (anonymous):

is it making sense or are you not seeing it yet?

OpenStudy (anonymous):

im seeing it in the answers theres A. -21x + 2 B. -21x - 14 C. -35x D. 21x + 2

OpenStudy (anonymous):

So its B

OpenStudy (anonymous):

because it cant be broken down anymore

OpenStudy (anonymous):

right.... that's correct, you've done all you can do. They didn't give you a value for x, so you can't do any more simplifying.

OpenStudy (anonymous):

6. Simplify: -6y + 4x + 2y - 10 + x

OpenStudy (anonymous):

Group the terms... all x terms together, all y terms together, and all numbers together

OpenStudy (anonymous):

-6y + 4x + 2y - 10 + x -6y + 2y + 4x + x - 10 -4y + 5x - 10

OpenStudy (anonymous):

Are you getting the idea?

OpenStudy (anonymous):

yes i got it ! thanks

OpenStudy (anonymous):

Glad to help :)